OpenStudy (anonymous):

arcsin (1/square root of 2)

OpenStudy (anonymous):

45 degrees and 135 degrees

OpenStudy (anonymous):

This means the angle whose sine is 1/sqrt 2. The domain of arcsin is 1,-1 so pi/4 and 3pi/4 (45 and 135 degrees).

OpenStudy (anonymous):

\[\sin^{-1} \left( 1\div \sqrt{2} \right)\] You mean that? Why doesn't anybody use LaTeX??

OpenStudy (anonymous):

\[Sin^{-1} \frac{1}{\sqrt{2}}\] Sorry, that better?